The line y= 2/3x +5 has a slope of m=2/3.What is the slope of a line perpendicular to this line?
Answer by ewatrrr(10682)


Hi
the slopes of perpendicular lines are negative reciprocals of each other.
line y= 2/3x +5 m = 2/3
m= -(3/2) would be the slope of a Line perpendicular to the above line